# L’approximation arithmétique de Fujita sur courbes adéliques

### 27 Jan 2024 v1

@@ Abstract
In this paper, we will prove an analogue of Fujita’s approximation theorem under the framework of Arakelov theory over adelic curves, which proves a conjecture of Huayi Chen and Atsushi Moriwaki.
@@
Dans cet article, on démontrera une analogie du théorème de l’approximation de Fujita sous le cadre de la théorie d’Arakelov sur courbes adéliques, qui démontre une conjecture de Huayi Chen et Atsushi Moriwaki.
@@ 1 Introduction
Fujita approximation is an approximative version of Zariski decomposition of pseudo-effective divisors [25]. More precisely, let $X$ be a projective variety over a field $k$ and $L$ be a line bundle on $X$. The volume of $L$ is defined as
$\operatorname{vol}(L)=\limsup_{D\rightarrow\infty}\frac{\dim_{k}H^{0}(X,L^{\otimes D})}{D^{\dim(X)}/(\dim(X))!}.$
We say that $L$ is a big line bundle if $\operatorname{vol}(L)>0$. Fujita’s approximation theorem states that, for all $\epsilon>0$, there exist a $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(L^{\otimes p})$ can be decomposed the tensor product of an ample line bundle $A$ and an effective line bundle $M$, which satisfy
@@
In fact, this result asserts that a big line bundle can be arbitrarily closely approximated by ample line bundles.
This approximated decomposition theorem is proved by T. Fujita [10] for the case of characteristic 00, and by S. Takagi [21] for an arbitrary characteristic. We refer the readers to [15, §11.4] for a survey on this topic.
@@ 1.1 Arithmetic analogue over number fields
It is quite nature to find an analogue of Fujita’s approximation theorem under the framework of Arakelov geometry, since we can define arithmetic positivities and arithmetic volume of a normed line bundle over arithmetic varieties.
@@
$\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes D})=\log\#\{s\in H^{0}(\mathscr{X},\mathscr{L}^{\otimes D})\mid\|s\|_{\sup,\sigma}\leqslant 1,\;\forall\sigma\in M_{K,\infty}\},$
@@
By the works mentioned above, the arithmetic volume function $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ has many similar behaviors to those of the geometric setting $\operatorname{vol}(\raisebox{1.72218pt}{.})$. So it is expected that an arithmetically big line bundle $\overline{\mathscr{L}}$, which satisfies $\widehat{\operatorname{vol}}(\overline{\mathscr{L}})>0$, also has the Fujita approximation. In fact, this used to be a conjecture of A. Moriwaki.
@@
In [2, 23], H. Chen and X. Yuan proved the arithmetic Fujita approximation over number field at almost the same time. Although they both applied [16, §3], but their methods are essential different. X. Yuan [23] takes pluri-subharmonic envelope to deal with the infinite part, which is well-known in complex analysis.
H. Chen applied his $\mathbb{R}$-filtration method developed in [3]. More precisely, by considering the successive minima on $E_{n}=H^{0}(\mathscr{X},\mathscr{L}^{\otimes n})$, we can make that $\bigoplus\limits_{n\geqslant 0}E_{n}$ is a filtrated graded linear series. Then by combining some properties of approximable graded linear series in [16], we apply some convergent properties of $\mathbb{R}$-filtrations to construct an ample line bundle which approximates $\overline{\mathscr{L}}$ well.
@@ 1.2 Arakelov geometry over adelic curves
In [7], H. Chen and A. Moriwaki developed a kind of Arakelov theory over so-called adelic curves. Based on this, they built the arithmetic intersection theory in [8], and studied the arithmetic positivities in [9]. Under this setting, we can study the arithmetic geometry over number fields, function fields and other bases with adelic structures by a united method.
@@
Briefly speaking, an adelic curve $S=(K,(\mathcal{A},\Omega,\nu),\phi)$ is an arbitrary field $K$ which is equipped with a parameter set $\Omega$ of all its absolute values. We require the set $\Omega$ parameterized by $\phi$ has a $\sigma$-algebra structure $\mathcal{A}$ and a measure $\nu$.
Under this setting, we can generalize many classic results and topics on Arakelov geometry, including the arithmetic volume function, arithmetic positivity, and Hilbert-Samuel formula. Since the adelic structure becomes more complicated, the obstructions original from measurability, integrability, and non-discreteness make the developing process more technique. In fact, some classic properties are not valid any longer in this framework.
@@
Inspired by [2, 23], it is quite natural to ask whether the Fujita approximation is still valid over adelic curves. Over a general adelic curve, we are lack of the Minkowskian property, so we cannot apply the filtrated method by successive minina. At the same time, some original constructions do not work any longer without the help of integral models.
In [9, §8.5], H. Chen and A. Moriwaki proved a similar result to the arithmetic Fujita approximation over adelic curves, which is called the relative Fujita approximation. More precisely, let an adelic line bundle $\overline{L}$ where $L$ is big on $X$. Then for any $t<\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$, there exists a $p\in\mathbb{N}$ sufficiently large and a birational morphism $\varphi:X^{\prime}\rightarrow X$, such that $(\varphi^{*}L)^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{N}$ is relatively ample and $\overline{M}$ is effective, and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})>pt$.
Compared with this relative version, the classic arithmetic Fujita approximation seems to take an "average" at every minima of minimal slope, and approximates all of them simultaneously. In fact, this is a conjecture proposed by Chen-Moriwaki in [9, Remark 9.2.7].
@@ 1.3 Main result
In this article, we prove the following result on the arithmetic Fujita approximation over adelic curves under the framework [7, 9]. More general than the conjecture proposed in [9, Remark 9.2.7], we prove the case of adelic $\mathbb{R}$-line bundle.
@@
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, which satisfies either $\mathcal{A}$ is discrete or has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be an arithmetically big adelic $\mathbb{R}$-line bundle on $X$ over $S$. Then for any $\epsilon>0$, there exist a $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(\overline{L}^{\otimes p})=\overline{A}\otimes\overline{M}$, where $\overline{A}$ is arithmetically ample and $\overline{M}$ is arithmetically effective. Moreover, the inequality
$p^{-(\dim(X)+1)}\widehat{\operatorname{vol}}(\overline{A})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon$
is verified.
@@
In [9], they develop a method which construct a graded linear series over the adelic curve of trivially valued field from the general setting by its Harder-Narasimhan filtration. In addition, we can modify the original norms by its spectral norm, which makes the graded linear series to be filtrated. Under this construction, the new graded linear series has some asymptotic invariants same as the original one. Then some ideas in [2, 16] can be absorbed in this work.
Once we prove the case of integral and $\mathbb{Q}$-line bundles, we can approximate an $\mathbb{R}$-line bundle by a $\mathbb{Q}$-line bundle in an effective direction, and then we can prove the case of $\mathbb{R}$-line bundle.
@@
In the application of $\mathbb{R}$-filtration method, we need to take some limits of some graded linear series and measures, and most operations only work when the objects are measurable and integrable. If either $\mathcal{A}$ is discrete or has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$, there will be no obstructions on the measurability or integrability, and see [7, §4] for more details of the necessity. Of course, in [2, 23], there is no such problem since the above condition is satisfied automatically.
@@ 1.4 Organization of article
In §2, we introduce the notions of Arakelov theory over adelic curves and $\mathbb{R}$-filtration. In §3, we consider graded linear series over a general adelic curve and construct those over the adelic curve of trivially valued field filtrated by the Harder-Narasimhan filtration. In §4, we prove the property of Fujita approximation of certain graded linear series. In §5, we prove the arithmetic Fujita approximation of adelic line bundles.
@@ Acknowledgements
I would like to thank Prof. Atsushi Moriwaki for introducing this subject to me. In addition, I would like to thank Prof. Huayi Chen for many useful discussions and suggestions, and Dr. Wenbin Luo for the advices on some key techniques. Parts of this work was accomplished during my visit to Institute for Theoretical Sciences, Westlake University, and I would like to thank their generous reception. Chunhui Liu was supported by National Science Foundation of China No. 12201153.
@@ 2 Arakelov theory over adelic curves
In this section, we will introduce the foundations of Arakelov over adelic curves and $\mathbb{R}$-filtration. The main references are [7, 9].
@@ 2.1 Foundations of Arakelov geometry over adelic curves
In this part, we recall some useful knowledge on the Arakelov geometry over adelic curves, see [7] for a systematical introduction to this subject.
@@ 2.1.1 Adelic curve
We say $(\Omega,\mathcal{A},\nu)$ an adelic structure on $K$ equipped with a mapping
@@
    is $\mathcal{A}$-measurable, integrable with respect to $\nu$.
@@
We denote by $\Omega_{f}$ the subset of $\Omega$ whose images in $M_{K}$ with respect to $\phi$ is non-archimedean, and by $\Omega_{\infty}$ the sbuset of $\Omega$ whose image in $M_{K}$ is archimedean. We have $\Omega=\Omega_{f}\sqcup\Omega_{\infty}$ by definition directly.
@@
We refer the readers to [7, §3] for more details on this notion.
@@
Let $K$ be a number field or a function field. Then the usual adelic ring $\mathbb{A}_{K}$ gives an adelic structure, and constructs examples of proper adelic curves.
Let $K=\mathbb{Q}(\varphi_{1},\ldots,\varphi_{d})$ be a finitely transcendental extension over $\mathbb{Q}$. In [17], A. Moriwaki constructed a special adelic structure over $K$, and defined a height function with respect to it. This height function has been applied to study many arithmetic geometry subjects.
@@
We note $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ the adelic curve which counts the absolute value |.$|\raisebox{1.72218pt}{.}|_{0}$ only once. This example will be applied later.
We refer the readers to [7, §3.2] for more examples of adelic curve.
@@ 2.1.2 Adelic vector bundle
An adelic bundle $\overline{G}=(G,(\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w})_{w\in\Omega})$ is said to be an adelic quotient bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $G$ is a quotient vector space of $E$, and .$\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w}$ is the quotient norm .$\|\raisebox{1.72218pt}{.}\|_{w}$ with respect to the canonical homomorphism $E\rightarrow G$ for each $w\in\Omega$. If $G$ is the quotient space of $E$ modulo $F$, and equip with the quotient metrics from $\overline{F}$, then we denote $\overline{G}=\overline{E}/\overline{F}=\overline{E/F}$.
For an adelic vector bundle $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, we define $\det(\overline{E})=(\det(E),(\|\raisebox{1.72218pt}{.}\|_{\det,w})_{w\in\Omega})$ as follows: the vector $K$-space $\det(E)=\bigwedge^{\dim(E)}(E)$ is the determinant space of $E$, and .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$ is the determinant norm on $\det(E_{w})$ of .$\|\raisebox{1.72218pt}{.}\|_{w}$ for each $w\in\Omega$. We refer the readers to [7, §1.2.3] for the details of the construction of .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$.
@@
In general, for an adelic vector bundle $\overline{E}$ over $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, we cannot make sure that all its subspace equipped with the subspace norms is still an adelic vector bundle, and its quotient spaces admit the same situation. However, if $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$, we will get adelic vector bundles from all the above constructions. See [7, Proposition 4.1.24] for more details.

### 17 Mar 2024 v2

@@ Abstract
In this paper, we will prove an analogue of Fujita’s approximation under the framework of Arakelov theory over adelic curves, which gives a positive response to a conjecture of Huayi Chen and Atsushi Moriwaki.
@@
Dans cet article, on démontrera une analogie de l’approximation de Fujita sous le cadre de la théorie d’Arakelov sur courbes adéliques, qui donne une réponse positive d’une conjecture de Huayi Chen et Atsushi Moriwaki.
@@ 1 Introduction
Fujita approximation is an approximative version of Zariski decomposition of pseudo-effective divisors, and see [28] for more details. More precisely, let $X$ be a projective variety over a field $k$, and $L$ be a line bundle on $X$. The volume of $L$ is defined as
$\operatorname{vol}(L)=\limsup_{D\rightarrow\infty}\frac{\dim_{k}H^{0}(X,L^{\otimes D})}{D^{\dim(X)}/\dim(X)!}.$
We say that $L$ is a big line bundle if $\operatorname{vol}(L)>0$. Fujita’s approximation theorem states that, for all $\epsilon>0$, there exist a $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(L^{\otimes p})$ can be decomposed as the tensor product of an ample line bundle $A$ and an effective line bundle $M$, which satisfy
@@
Intuitively, this result asserts that a big line bundle can be arbitrarily closely approximated by ample line bundles along an effective direction.
This approximated decomposition theorem is proved by T. Fujita [12] for the case of characteristic 00, and by S. Takagi [24] for an arbitrary characteristic. We refer the readers to [17, §11.4] for a survey on this topic.
@@ 1.1 Arithmetic analogue over number fields
It is quite nature to ask whether there is an analogue of Fujita’s approximation theorem under the framework of Arakelov geometry, since we can define the arithmetic positivity and arithmetic volume of an adelic line bundle over arithmetic varieties.
@@
$\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes D})=\log\#\left\{s\in H^{0}(\mathscr{X},\mathscr{L}^{\otimes D})\mid\|s\|_{\sup,v}\leqslant 1,\;\forall\;v\in M_{K,\infty}\right\},$
@@
By the works mentioned above, the arithmetic volume function $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ has many similar properties to those of the geometric setting $\operatorname{vol}(\raisebox{1.72218pt}{.})$, so it is expected that an arithmetically big line bundle $\overline{\mathscr{L}}$, which means $\widehat{\operatorname{vol}}(\overline{\mathscr{L}})>0$, also has the property of Fujita approximation. In fact, this used to be a conjecture of A. Moriwaki.
@@
In [4, 26], H. Chen and X. Yuan proved the arithmetic Fujita approximation over number field at almost the same time. Although they both applied [18, §3], but their methods are essential different. X. Yuan [26, Theorem C] takes pluri-subharmonic envelope to deal with the infinite part, which is well-known in complex analysis.
In [4], H. Chen applied his $\mathbb{R}$-filtration method developed in [5]. More precisely, by considering the successive minima on $H^{0}(\mathscr{X},\mathscr{L}^{\otimes n})$, we can make $\bigoplus\limits_{n\geqslant 0}H^{0}(\mathscr{X},\mathscr{L}^{\otimes n})$ to be a filtrated graded linear series. Then by combining some results of approximable graded linear series in [18], we are able to apply some convergent properties of $\mathbb{R}$-filtrations to construct an ample line bundle which approximates $\overline{\mathscr{L}}$ well. This is the main idea of the proof in [4].
@@ 1.2 Arakelov geometry over adelic curves
In [9], H. Chen and A. Moriwaki developed a kind of Arakelov theory over the so-called adelic curve. Based on this framework, they built an arithmetic intersection theory in [10], and studied the arithmetic positivities in [11]. Under this setting, we can study the arithmetic geometry over number fields, function fields and other bases with adelic structures under a same framework.
@@
Briefly speaking, an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ consists of an arbitrary field $K$ which is equipped with a parameter set $\Omega$ of all its absolute values. We require the set $\Omega$ parameterized by $\phi$ has a $\sigma$-algebra structure $\mathcal{A}$ and a measure $\nu$ compatible with $\mathcal{A}$.
Under this setting, we can generalize many classic results and subjects of Arakelov geometry, including the arithmetic volume function, arithmetic positivity, and Hilbert-Samuel formula. Since the adelic structure becomes more complicated than those of classic cases, the obstructions original from measurability, integrability, and non-discreteness make the developing process more technique. In fact, some classic properties are not valid any longer in this framework, though we have developed many in [9, 10, 11], and see also [23].
@@
Inspired by [4, 26], it is quite natural to ask whether the Fujita approximation is still valid over adelic curves. Over a general adelic curve, we are lack of the Minkowskian property, so we cannot apply the filtrated method by successive minina directly in [4] At the same time, some constructions do not work any longer without the help of integral models.
In [11, §8.5], H. Chen and A. Moriwaki proved a similar result to the arithmetic Fujita approximation over adelic curves, which is called the relative Fujita approximation. More precisely, let $\overline{L}$ be an adelic line bundle, where $L$ is big on $X$. Then for any $t<\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$, there exists a sufficiently large $p\in\mathbb{N}$ and a birational morphism $\varphi:X^{\prime}\rightarrow X$, such that $(\varphi^{*}\overline{L})^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{N}$ is relatively ample and $\overline{M}$ is effective, and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})>pt$. Here $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ denote the asymptotic maximal and minimal slopes respectively.
Compared with this relative version above, the classic arithmetic Fujita approximation seems to take an "average" at every minima of minimal slope, and approximates all of them simultaneously, not only the maximal one. In fact, this is conjectured by Chen-Moriwaki in [11, Remark 9.2.7].
@@ 1.3 Main result
In this article, we prove the following result on the arithmetic Fujita approximation over adelic curves under the framework [9, 11], which gives a positive response to the conjecture [11, Remark 9.2.7]. More generally, we prove a version of adelic $\mathbb{R}$-line bundle below at the same time, which is also a generalization of [4, Theorem 4.3] and [26, Theorem C].
@@
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, which satisfies either $\mathcal{A}$ is discrete or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be an arithmetically big $\mathbb{R}$-line bundle on $X$ over $S$. Then for any $\epsilon>0$, there exist a sufficiently large $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(\overline{L}^{\otimes p})=\overline{A}\otimes\overline{M}$, where $\overline{A}$ is arithmetically ample and $\overline{M}$ is an arithmetically effective $\mathbb{R}$-line bundle. Moreover, we have
$p^{-(\dim(X)+1)}\widehat{\operatorname{vol}}(\overline{A})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon.$
@@
In [11, §3], they develop a method which constructs a graded linear series over the adelic curve of trivially valued field from general one by its Harder-Narasimhan filtration. Next, we can modify the original norms by its spectral norm, which makes the graded linear series to be filtrated. Under this construction, the new graded linear series has some asymptotic invariants same as the original one. Then some ideas in [4, 18] can be absorbed in this work.
Once we prove the case of $\mathbb{Q}$-line bundle (in Corollary 5.2.1) from the case of usual adelic line bundle (in Corollary 5.1), we can approximate an arithmetically big $\mathbb{R}$-line bundle by an arithmetically big $\mathbb{Q}$-line bundle in an effective direction, and then we can prove the case of $\mathbb{R}$-line bundle (in Theorem 5.2.2).
@@
In the application of $\mathbb{R}$-filtration method, we need to take some limits of some graded linear series and measures, and most operations only work when the objects are measurable and integrable. If either $\mathcal{A}$ is discrete or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$, there will be no obstructions on the measurability or integrability, and see [9, §4] for more details of the necessity.
In order to study some arithmetic positivities by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$, we need to suppose the base field $K$ is perfect. For example, for an adelic line bundle $\overline{L}$ where $L$ is nef, we need to apply $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})\geqslant 0$ to deduce $\overline{L}$ is arithmetically nef by [11, Proposition 9.1.7], where we need the perfect condition. See [11, §9] for more details of the necessity.
Of course, in [4, 26], there is no such problem since for the case of number field, because $\mathcal{A}$ is discrete in the language of adelic curve and the base field is perfect by its characteristic 00.
@@ 1.4 Organization of article
In §2, we introduce the required notions of Arakelov theory over adelic curves and $\mathbb{R}$-filtration. In §3, we consider graded linear series over a general adelic curve and construct those over the adelic curve of trivially valued field filtrated by the Harder-Narasimhan filtration. In §4, we prove the property of Fujita approximation of certain graded linear series. In §5, we prove the arithmetic Fujita approximation of adelic line bundles in the usual case, the case of $\mathbb{Q}$-line bundle and the case of $\mathbb{R}$-line bundle sequentially.
@@ Acknowledgements
I would like to thank Prof. Atsushi Moriwaki for introducing this subject to me. In addition, I would like to thank Prof. Huayi Chen for many useful discussions and suggestions, and Dr. Wenbin Luo for the advices on some key techniques. Part of this work was accomplished during my visit to Institute for Theoretical Sciences, Westlake University, and I would like to thank their generous invitation and reception. Chunhui Liu was supported by National Science Foundation of China No. 12201153.
@@ 2 Arakelov theory over adelic curves
In this section, we will introduce the foundations of Arakelov theory over adelic curves and of $\mathbb{R}$-filtration. The main references are [9, 11].
@@ 2.1 Foundations of Arakelov geometry over adelic curves
First, we recall some useful knowledge on the Arakelov geometry over adelic curves, and see [9] for a systematical introduction to this subject.
@@ 2.1.1 Adelic curve
We say that $(\Omega,\mathcal{A},\nu)$ is an adelic structure on $K$ equipped with a mapping
@@
    is $\mathcal{A}$-measurable and integrable with respect to $\nu$.
@@
We denote by $\Omega_{f}$ the subset of $\Omega$ whose images in $M_{K}$ with respect to $\phi$ is non-archimedean, and by $\Omega_{\infty}$ the subset of $\Omega$ whose image in $M_{K}$ is archimedean. We have $\Omega=\Omega_{f}\sqcup\Omega_{\infty}$ by definition directly.
@@
We refer the readers to [9, §3] for more details on these notions.
@@
Let $K$ be a number field or a function field. Then the usual adèle ring of $K$ gives an adelic structure, and constructs examples of proper adelic curves.
Let $K$ be a finitely transcendental extension over $\mathbb{Q}$, which is a sub-field of $\mathbb{C}$. In [19], A. Moriwaki constructed a special adelic structure over $K$, and defined a height function with respect to it. This height function has been applied to study many arithmetic geometry subjects.
@@
We note $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ the adelic curve which counts the absolute value |.$|\raisebox{1.72218pt}{.}|_{0}$ once. By definition, $S_{0}$ is a proper adelic curve.
Example 2.1.1 will be used in the article later. We refer the readers to [9, §3.2] for more examples of adelic curve.
@@ 2.1.2 Adelic vector bundle
An adelic bundle $\overline{G}=(G,(\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w})_{w\in\Omega})$ is said to be an adelic quotient bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $G$ is a quotient vector space of $E$, and .$\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w}$ is the quotient norm .$\|\raisebox{1.72218pt}{.}\|_{w}$ with respect to the canonical quotient homomorphism $E\rightarrow G$ for each $w\in\Omega$. If $G$ is the quotient space of $E$ modulo $F$, and is equipped with the quotient metrics from $\overline{F}$, then we denote $\overline{G}=\overline{E}/\overline{F}=\overline{E/F}$.
For an adelic vector bundle $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, we define $\det(\overline{E})=(\det(E),(\|\raisebox{1.72218pt}{.}\|_{\det,w})_{w\in\Omega})$ as follows: the $K$-vector space $\det(E)=\bigwedge^{\dim(E)}(E)$ is the determinant space of $E$, and .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$ is the determinant norm on $\det(E)_{w}$ of .$\|\raisebox{1.72218pt}{.}\|_{w}$ for each $w\in\Omega$. We refer the readers to [9, §1.2.3] for the details of the construction of .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$.
@@
In general, for an adelic vector bundle $\overline{E}$ over $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, we cannot make sure that all its subspace equipped with the subspace norms is still an adelic vector bundle, and its quotient spaces admit the same situation. However, if $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$, we will get adelic vector bundles from all the above constructions, and see [9, Proposition 4.1.24] for more details. In the whole article, we will always assume these conditions.